# Heterogeneity-Oblivious Robust Federated Learning

### v1

@@ III Motivation
However, as a plug-in, LoRA modifies a layer (base weight is $W$) by a low-rank increment without altering the layer’s original input-output mapping, e.g., $W^{\prime}=W+\Delta W,\Delta W=BA$, $A\in\mathbb{R}^{r\times d_{\text{in}}},B\in\mathbb{R}^{d_{\text{out}}\times r}$, where $A$ & $B$ denote LoRA-A and LoRA-B, respectively, and $r\ll\min(d_{\text{in}},d_{\text{out}})$, $d_{\text{in}}$ and $d_{\text{out}}$ are the input and output dimensions of the injected layer. As a result, LoRA inherits heterogeneity-induced inconsistencies. Naively averaging LoRA updates can remain unstable under hyper-heterogeneity. We therefore conduct extensive analyses to identify the stable components across clients, which in turn guide our heterogeneity-oblivious detection and robust aggregation design. Specifically, we use 10 clients spanning two architecture (CNNs and RNNs), and evaluate on CIFAR-10/100 and FMNIST (partitioned non-IID via different Dirichlet sampler), and analyze client behavior under benign and six representative poisoning attacks. More experimental details appear in § VI. Notably, figures in this section show one CIFAR10-LIE-$\alpha{=}0.5$ instance due to space, but the same qualitative trends hold across datasets, $\alpha$, architectures, and attacks. We summarize the key observations below.
@@
To evaluate the stability of LoRA under hyper-heterogeneity and poisoning attack, we functionally decouple LoRA as $A$ & $B$, and perform singular value decomposition (SVD) on them. Then, Top-$k$ singular value energy ratio is used to analyze the distribution of singular values in the $A$ and $B$ across different model layers and clients [9]. This metric measures how much an update concentrates in $k$ dominant directions. A high ratio means a small set of principal components carries most of the signal, making it easier to detect poisoning that break this directional pattern, whereas a low ratio indicates noisy updates that are more prone to heterogeneity or poisoning. The choice of $k$ follows the paper [9].
@@ IV Preliminary
Federated learning enables distributed clients to collaboratively train without exposing raw data. To address the robustness challenges brought by the hyper-heterogeneity, according to our observation, we insert trainable low-rank adaptations into specific stable layers (i.e., feature-first and classifier). In LoRA, given a original weight $W$, its update $\Delta W$ is confined to a structured subspace by $\Delta W=BA$, where $A$ and $B$ denote LoRA-A and LoRA-B, respectively. Each client retains a private backbone, and only LoRAs are shared during training, which is free from the communication restriction and shrinks the attack surface.
For $N$ participating clients, after local training, each client $c$ retains their private model backbone, and only shares $(A_{c},B_{c})$. The server uses a detection algorithm to kick out the poisoned clients, and applies an aggregation rule $\mathcal{A}(\cdot)$ to compute:
@@
Empirically, we find that LoRA-A exhibits stronger stability, while LoRA-B is more sensitive to heterogeneity. Hence, we functionally decouple them: $A$ is used for poisoning detection, while $A$ & $B$ support aggregation. This separation enables effective detection and aggregation under hyper-heterogeneity.
@@ V-C1 Dimensional Alignment
For client $c$, direct element-wise aggregation is infeasible the sizes of $A_{c}\!\in\!\mathbb{R}^{r\times d_{\text{in}}^{(c)}}$ and $B_{c}\!\in\!\mathbb{R}^{d_{\text{out}}^{(c)}\times r}$ are layer-specific. Thus, we zero-pad each matrix to the global maximum shape: $\tilde{A}_{c}\in\mathbb{R}^{r\times d_{\text{in}}^{\max}}$, $\qquad\tilde{B}_{c}\in\mathbb{R}^{d_{\text{out}}^{\max}\times r}$, where $d_{\text{in}}^{\max}$ and $d_{\text{out}}^{\max}$ are the maximum input/output dimensions across clients. However, the padded regions are meaningless, and direct averaging padded LoRA dilutes the meaningful updates. To prevent zero-padding from skewing the result, we introduce binary masks $M_{c}^{A},M_{c}^{B}\in\{0,1\}^{\text{matching shape}}$, with $1$ on valid entries and 0 on padded entries. Masked averaging is then performed element-wise:

### v2

@@ III Motivation
However, as a plug-in, LoRA modifies a layer (base weight is $W$) by a low-rank increment without altering the layer’s original input-output mapping, e.g., $W^{\prime}=W+\Delta W,\Delta W=BA$, $A\in\mathbb{R}^{r\times d_{\text{in}}},B\in\mathbb{R}^{d_{\text{out}}\times r}$, where AA & $B$ denote LoRA-A and LoRA-B, respectively, and $r\ll\min(d_{\text{in}},d_{\text{out}})$, $d_{\text{in}}$ and $d_{\text{out}}$ are the input and output dimensions of the injected layer. As a result, LoRA inherits heterogeneity-induced inconsistencies. Naively averaging LoRA updates can remain unstable under hyper-heterogeneity. We therefore conduct extensive analyses to identify the stable components across clients, which in turn guide our heterogeneity-oblivious detection and robust aggregation design. Specifically, we use 10 clients spanning two architecture (CNNs and RNNs), and evaluate on CIFAR-10/100 and FMNIST (partitioned non-IID via different Dirichlet sampler), and analyze client behavior under benign and six representative poisoning attacks. More experimental details appear in § VI. Notably, figures in this section show one CIFAR10-LIE-$\alpha{=}0.5$ instance due to space, but the same qualitative trends hold across datasets, $\alpha$, architectures, and attacks. We summarize the key observations below.
@@
To evaluate the stability of LoRA under hyper-heterogeneity and poisoning attack, we functionally decouple LoRA as AA & $B$, and perform singular value decomposition (SVD) on them. Then, Top-$k$ singular value energy ratio is used to analyze the distribution of singular values in the AA and $B$ across different model layers and clients [9]. This metric measures how much an update concentrates in $k$ dominant directions. A high ratio means a small set of principal components carries most of the signal, making it easier to detect poisoning that break this directional pattern, whereas a low ratio indicates noisy updates that are more prone to heterogeneity or poisoning. The choice of $k$ follows the paper [9].
@@ IV Preliminary
Federated learning enables distributed clients to collaboratively train without exposing raw data. To address the robustness challenges brought by the hyper-heterogeneity, according to our observation, we insert trainable low-rank adaptations into specific stable layers (i.e., feature-first and classifier). In LoRA, given a original weight $W$, its update $\Delta W$ is confined to a structured subspace by $\Delta W=BA$, where AA and $B$ denote LoRA-A and LoRA-B, respectively. Each client retains a private backbone, and only LoRAs are shared during training, which is free from the communication restriction and shrinks the attack surface.
For NN participating clients, after local training, each client $c$ retains their private model backbone, and only shares $(A_{c},B_{c})$. The server uses a detection algorithm to kick out the poisoned clients, and applies an aggregation rule $\mathcal{A}(\cdot)$ to compute:
@@
Empirically, we find that LoRA-A exhibits stronger stability, while LoRA-B is more sensitive to heterogeneity. Hence, we functionally decouple them: AA is used for poisoning detection, while AA & $B$ support aggregation. This separation enables effective detection and aggregation under hyper-heterogeneity.
@@ V-C1 Dimensional Alignment
For client $c$, direct element-wise aggregation is infeasible the sizes of $A_{c}\!\in\!\mathbb{R}^{r\times d_{\text{in}}^{(c)}}$ and $B_{c}\!\in\!\mathbb{R}^{d_{\text{out}}^{(c)}\times r}$ are layer-specific. Thus, we zero-pad each matrix to the global maximum shape: $\tilde{A}_{c}\in\mathbb{R}^{r\times d_{\text{in}}^{\max}}$, $\qquad\tilde{B}_{c}\in\mathbb{R}^{d_{\text{out}}^{\max}\times r}$, where $d_{\text{in}}^{\max}$ and $d_{\text{out}}^{\max}$ are the maximum input/output dimensions across clients. However, the padded regions are meaningless, and direct averaging padded LoRA dilutes the meaningful updates. To prevent zero-padding from skewing the result, we introduce binary masks $M_{c}^{A},M_{c}^{B}\in\{0,1\}^{\text{matching shape}}$, with 11 on valid entries and 0 on padded entries. Masked averaging is then performed element-wise: